# History Matters: Temporal Knowledge Editing in Large Language Model

### 09 Dec 2023 v1

@@ 1 Introduction
Our code and dataset will be released to the community.
2 Temporal Knowledge Editing
@@
2.1 Background: Knowledge in Language Models
@@
2.2 Requirements for Knowledge Editing
@@ 2.3 Task Definition
As mentioned in Section 2.1, we use triples to represent knowledge. In addition, we attach temporal scope to each relational fact $\{(s,r,o,t_{s},t_{u})\}$ since the relational fact often shows temporal dynamics, where $t_{s}$ and $tu$ denote the fact $(s,r,o)$ valid from $ts$ to $tu$. As time passes, the same subject and relation will correspond to different objects in sequence, such as (U.S., president, Donald Trump, 2017, 2021), (U.S., president, Joseph Biden, 2021, N/A) and so on. Consistent with previous work, we use $(s,r,o,t_{s},t_{u})\rightarrow(s,r,o^{\star},t_{s}^{\star},t_{u}^{\star})$ to represent an knowledge editing operation, denoted as $e$, meaning that the object of subject $s$ under relation $r$ changes from $o$ to $o^{\star}$. And the reason for the edit is that $(s,r,o)$ is valid from $t_{s}$ until $t_{u}$, and after that $(s,r,o^{\star})$ is valid from $t_{s}^{\star}$ to $t_{u}^{\star}$ [^2] Note that the $t_{u}$ is vacant if the knowledge is still valid currently.. Specifically, $o$ and $o^{\star}$ can be the same object, at which point the editing operation $(s,r,o,t_{s},t_{u})\rightarrow(s,r,o,t_{s},t_{u}^{\star})$ will make the model know that the fact $(s,r,o)$ lasts until $t_{u}^{\star}$, where $t_{u}^{\star}$ is later than $t_{u}$. For convenience of expression, we refer to $(s,r,o,t_{s},t_{u})$ as historical knowledge and $(s,r,o^{\star},t_{s}^{\star},t_{u}^{\star})$ as current knowledge.
@@
Given a collection of knowledge editing operations =$\mathcal{E}=\{e_{1},e_{2},\dots\}$ and a language model $\mathcal{M}$, knowledge editor $F$ aims to get an edited model $\mathcal{M}_{e}=F(\mathcal{M},\mathcal{E})$ that satisfy the requirements mentioned in Section 2.2. Specifically, in temporal knowledge editing, for the edited model $\mathcal{M}$, we expect that it should not only memorize the current knowledge $(s,r,o^{\star})$, but the knowledge in it should also be organized on the timeline. In other words, the model can answer correctly when asked questions about the “previous” or “last one” knowledge $(s,r,o)$, or when asked directly about the relevant knowledge $(s,r,?)$ at a certain moment in time between $t_{s}$ and $t_{u}$. For example, after updating the model’s knowledge of U.S. president to Biden, the test questions include “Who was the former president of the United States?” or “Who was the president of the United States in 2005?” .
2.4 Evaluation of Temporal Knowledge Editing
@@
Since time is constantly moving forward, a piece of knowledge may be updated multiple times in different temporal scopes. Therefore, we define three evaluation subtasks: 1) Temporal Knowledge Single Editing (TSE), which requires that, after a single update to a piece of knowledge, the model sustains temporal chronology. 2) Temporal Knowledge Multiple Editing (TME), demanding ordered temporal succession after multiple unidirectional knowledge updates. It is worth pointing out the difference between TME and mass-editing of prior work (Meng et al. 2022b; Mitchell et al. 2022a): mass-editing refers to the simultaneous editing of multiple facts, which are usually unrelated, while TME refers to the continuous editing of the same $s$ and $r$ at different time scopes. 3) Temporal Knowledge Extending Editing (TEE), which requires that duration of target knowledge perceived by the model is successfully prolonged. More details on the benchmark are presented in Section 3.
3 AToKe: Assessment of Temporal Knowledge Editing
@@
3.1 Basic Temporal Knowledge Collection
@@
Our dataset is based on YAGO3.0.3333https://yago-knowledge.org/downloads/yago-3 (Mahdisoltani, Biega, and Suchanek 2015), a knowledge base comprising fact triples associated with millions of entities extracted from Wikipedia, enriched with WordNet, GeoNames, and other data sources. Following the previous work (Dasgupta, Ray, and Talukdar 2018), we begin by sampling temporal facts from the YAGO data. We extract all fact triples with their time scope and obtain the temporal facts set $\{(s,r,o,t_{s},t_{u})\}$. Subsequently, after matching by $s$ and $r$ and sorting by $t_{s}$ with all these facts, we obtain the chain of temporal facts $\mathcal{C}=\{(s,r,o_{1},t_{s_{1}},t_{u_{1}}),...,(s,r,o_{N},t_{s_{N}},t_{u_{N}})\}$. Finally, to ensure non-overlapping and non-contradictory time sequences within the chain, we employ a heuristic algorithm. See the Appendix A.1 for more technical details.
@@
For a chain of temporal facts $\{(s,r,o_{N},t_{s_{N}},t_{u_{N}})\}$, its $o$, $t_{s}$ and $t_{u}$ are constantly changing. Therefore, to test the performance of knowledge editing, we need to determine where the knowledge is located in the model. First, we use the GPT-J model [^4] Without loss of generality, we use GPT-J as the representative LLM used in the experiments. to verify and filter out $(s,r,*)$ that the model does not have. Next, we locate the model’s knowledge position in the temporal fact chain. By asking questions about the facts in the temporal chain, we find the most recent fact in each chain that the model has successfully recalled as the model’s initial knowledge. See Appendix A.1 for more details of locating.
@@
As mentioned in Section 2.1, given a chain of temporal facts that we construct, we construct template $q_{r}(\cdot)$ for each relation $r$. The constructed questions are categorized into two categories based on time: 1) Explicit time question, which refers to asking the question that have an explicit time frame. 2) Relative time question, on the other hand, does not contain an explicit time frame and uses words like “present”, “previous” or “last one” instead. We first utilize ChatGPT (gpt-3.5-turbo) to automatically generate questions for all relations, and then manually filter them. The details of question generation such as prompts used, question examples and statistics of questions are shown in Appendix A.1.
3.2 Construction of Three Datasets
@@ 3.3 Dataset Summary
The three datasets have a similar data format, but they differ in the number of edits and the questions used to evaluation. These datasets all contain edit sets (SE&ME&EE), questions ($\mathcal{Q}$), standardized answers to questions ($\mathcal{A}$), and answer alias sets. In these datasets, both single edit and multiple edits contain questions about historical and current knowledge, while extending edits contain only questions about the current knowledge. Single edit and extending edit have only one editing operation and corresponding questions for one $(s,r)$, while multiple edits contain multiple operations and questions. Format examples for the three datasets can be found in the Appendix A.2
@@
In the base temporal data collection phase, we first sample 169,996 temporal fact chains. After filtering by GPT-J, there are 13 different relations left and the number of our temporal chains is reduced to 8,820, which is the final number of the three datasets. For the length distribution of temporal fact chains, fact chains of length 2-5 account for more than 90%, with very few chains exceeding 10. In addition, the relationship ”playsFor” has the largest number of fact chains. The time is measured in years. And for the time scope, before adding the future fake facts, the latest knowledge is obtained in 2022, and after that it is extended to 2028. More detailed statistics are in the Appendix A.2.
3.4 Evaluation Metrics
@@
In particular, in AToKe-ME, we treat each edit as a single edit and compute the score as in AToKe-SE, and finally report the average of all edits. In addition, we calculate HES for all the historical facts at the end of multiple consecutive edits of the model to measure the overall editing performance (denoted as $*{}^{*}$). In AToKe-EE, since there is no historical knowledge, we do not need to measure HRS and HES. CThe results of detailed formal definitions and another metric can be found in the Appendix A.3.
4 Current Status on TKE
@@ 4.1 Experimental Setup
All of the above model editing methods make changes to the parameters of the model, which is the focus of our exploration. In addition, there are some methods based on external storage that do not edit the model. The performance of that class of methods depends on the knowledge classifiers or retrievers, which is analyzed and explored in the Appendix B.
@@ 4.2 Results and Analysis
5 METO: Multi-Editing with Time Objective
As we can see from the previous experiments in Section 4.2, the existing methods perform excellently in memorizing new knowledge, but catastrophic forgetting of historical knowledge occurs after editing. To alleviate this serious problem, we propose a simple editing framework METO (Multi-Editing with Time Objective) that can be applied to enhance existing editing methods easily.
@@ 5.1 Methodology
For a chain of temporal facts $\mathcal{C}=\{(s,r,o_{1},t_{s_{1}},t_{u_{1}}),...,(s,r,o_{N},t_{s_{N}},t_{u_{N}})\}$, the object is constantly changing and the knowledge of the model about $(s,r,\cdot)$ is determined by the collection time of its training corpus. Therefore, in order to better preserve the historical knowledge inside the model, we need to extract the model-time knowledge about $(s,r,\cdot)$. With the method in Section 3.1, we obtain the fact of what is happening under the model time which is noted as $C_{m}=\{(s,r,o_{i},t_{s_{i}},t_{u_{i}})\}$. By comparing $\mathcal{C}$ and $C_{m}$, we can get the knowledge that needs to be newly captured by the model $C_{m+}=\{(s,r,o_{i+1},t_{s_{i+1}},t_{u_{i+1}}),...,(s,r,o_{N},t_{s_{N}},t_{u_{N}})\}$.
@@
5.2 Results on AToKe-SE and AToKe-ME

### 14 Dec 2023 v3

@@ Introduction
Our benchmark has been released to the community to facilitate future research  [^1] https://github.com/Arvid-pku/ATOKE.
Temporal Knowledge Editing
@@
Background: Knowledge in Language Models
@@
Requirements for Knowledge Editing
@@
As mentioned in Section Background: Knowledge in Language Models, we use triples to represent knowledge. In addition, we attach temporal scope to each relational fact $\{(s,r,o,t_{s},t_{u})\}$ since the relational fact often shows temporal dynamics, where $t_{s}$ and $tu$ denote the fact $(s,r,o)$ valid from $ts$ to $tu$. As time passes, the same subject and relation will correspond to different objects in sequence, such as (U.S., president, Donald Trump, 2017, 2021), (U.S., president, Joseph Biden, 2021, N/A) and so on. Consistent with previous work, we use $(s,r,o,t_{s},t_{u})\rightarrow(s,r,o^{\star},t_{s}^{\star},t_{u}^{\star})$ to represent an knowledge editing operation, denoted as $e$, meaning that the object of subject $s$ under relation $r$ changes from $o$ to $o^{\star}$. And the reason for the edit is that $(s,r,o)$ is valid from $t_{s}$ until $t_{u}$, and after that $(s,r,o^{\star})$ is valid from $t_{s}^{\star}$ to $t_{u}^{\star}$ [^3] Note that the $t_{u}$ is vacant if the knowledge is still valid currently.. Specifically, $o$ and $o^{\star}$ can be the same object, at which point the editing operation $(s,r,o,t_{s},t_{u})\rightarrow(s,r,o,t_{s},t_{u}^{\star})$ will make the model know that the fact $(s,r,o)$ lasts until $t_{u}^{\star}$, where $t_{u}^{\star}$ is later than $t_{u}$. For convenience of expression, we refer to $(s,r,o,t_{s},t_{u})$ as historical knowledge and $(s,r,o^{\star},t_{s}^{\star},t_{u}^{\star})$ as current knowledge.
@@
Given a collection of knowledge editing operations =$\mathcal{E}=\{e_{1},e_{2},\dots\}$ and a language model $\mathcal{M}$, knowledge editor $F$ aims to get an edited model $\mathcal{M}_{e}=F(\mathcal{M},\mathcal{E})$ that satisfy the requirements mentioned in Section Requirements for Knowledge Editing. Specifically, in temporal knowledge editing, for the edited model $\mathcal{M}$, we expect that it should not only memorize the current knowledge $(s,r,o^{\star})$, but the knowledge in it should also be organized on the timeline. In other words, the model can answer correctly when asked questions about the “previous” or “last one” knowledge $(s,r,o)$, or when asked directly about the relevant knowledge $(s,r,?)$ at a certain moment in time between $t_{s}$ and $t_{u}$. For example, after updating the model’s knowledge of U.S. president to Biden, the test questions include “Who was the former president of the United States?” or “Who was the president of the United States in 2005?” .
Evaluation of Temporal Knowledge Editing
@@
Since time is constantly moving forward, a piece of knowledge may be updated multiple times in different temporal scopes. Therefore, we define three evaluation subtasks: 1) Temporal Knowledge Single Editing (TSE), which requires that, after a single update to a piece of knowledge, the model sustains temporal chronology. 2) Temporal Knowledge Multiple Editing (TME), demanding ordered temporal succession after multiple unidirectional knowledge updates. It is worth pointing out the difference between TME and mass-editing of prior work (Meng et al. 2022b; Mitchell et al. 2022a): mass-editing refers to the simultaneous editing of multiple facts, which are usually unrelated, while TME refers to the continuous editing of the same $s$ and $r$ at different time scopes. 3) Temporal Knowledge Extending Editing (TEE), which requires that duration of target knowledge perceived by the model is successfully prolonged. More details on the benchmark are presented in Section AToKe: Assessment of Temporal Knowledge Editing.
AToKe: Assessment of Temporal Knowledge Editing
@@
Basic Temporal Knowledge Collection
@@
Our dataset is based on YAGO3.0.3444https://yago-knowledge.org/downloads/yago-3 (Mahdisoltani, Biega, and Suchanek 2015), a knowledge base comprising fact triples associated with millions of entities extracted from Wikipedia, enriched with WordNet, GeoNames, and other data sources. Following the previous work (Dasgupta, Ray, and Talukdar 2018), we begin by sampling temporal facts from the YAGO data. We extract all fact triples with their time scope and obtain the temporal facts set $\{(s,r,o,t_{s},t_{u})\}$. Subsequently, after matching by $s$ and $r$ and sorting by $t_{s}$ with all these facts, we obtain the chain of temporal facts $\mathcal{C}=\{(s,r,o_{1},t_{s_{1}},t_{u_{1}}),...,(s,r,o_{N},t_{s_{N}},t_{u_{N}})\}$. Finally, to ensure non-overlapping and non-contradictory time sequences within the chain, we employ a heuristic algorithm.
@@
For a chain of temporal facts $\{(s,r,o_{N},t_{s_{N}},t_{u_{N}})\}$, its $o$, $t_{s}$ and $t_{u}$ are constantly changing. Therefore, to test the performance of knowledge editing, we need to determine where the knowledge is located in the model. First, we use the GPT-J model [^5] Without loss of generality, we use GPT-J as the representative LLM used in the experiments. to verify and filter out $(s,r,*)$ that the model does not have. Next, we locate the model’s knowledge position in the temporal fact chain. By asking questions about the facts in the temporal chain, we find the most recent fact in each chain that the model has successfully recalled as the model’s initial knowledge.
@@
As mentioned in Section Background: Knowledge in Language Models, given a chain of temporal facts that we construct, we construct template $q_{r}(\cdot)$ for each relation $r$. The constructed questions are categorized into two categories based on time: 1) Explicit time question, which refers to asking the question that have an explicit time frame. 2) Relative time question, on the other hand, does not contain an explicit time frame and uses words like “present”, “previous” or “last one” instead. We first utilize ChatGPT (gpt-3.5-turbo) to automatically generate questions for all relations, and then manually filter them.
Construction of Three Datasets
@@
The three datasets have a similar data format, but they differ in the number of edits and the questions used to evaluation. These datasets all contain edit sets (SE&ME&EE), questions ($\mathcal{Q}$), standardized answers to questions ($\mathcal{A}$), and answer alias sets. In these datasets, both single edit and multiple edits contain questions about historical and current knowledge, while extending edits contain only questions about the current knowledge. Single edit and extending edit have only one editing operation and corresponding questions for one $(s,r)$, while multiple edits contain multiple operations and questions.
@@
In the base temporal data collection phase, we first sample 169,996 temporal fact chains. After filtering by GPT-J, there are 13 different relations left and the number of our temporal chains is reduced to 8,820, which is the final number of the three datasets. For the length distribution of temporal fact chains, fact chains of length 2-5 account for more than 90%, with very few chains exceeding 10. In addition, the relationship ”playsFor” has the largest number of fact chains. The time is measured in years. And for the time scope, before adding the future fake facts, the latest knowledge is obtained in 2022, and after that it is extended to 2028.
Evaluation Metrics
@@
In particular, in AToKe-ME, we treat each edit as a single edit and compute the score as in AToKe-SE, and finally report the average of all edits. In addition, we calculate HES for all the historical facts at the end of multiple consecutive edits of the model to measure the overall editing performance (denoted as $*{}^{*}$). In AToKe-EE, since there is no historical knowledge, we do not need to measure HRS and HES.
Current Status on TKE
@@
All of the above model editing methods make changes to the parameters of the model, which is the focus of our exploration. In addition, there are some methods based on external storage that do not edit the model. The performance of that class of methods depends on the knowledge classifiers or retrievers.
@@
METO: Multi-Editing with Time Objective
As we can see from the previous experiments in Section Results and Analysis, the existing methods perform excellently in memorizing new knowledge, but catastrophic forgetting of historical knowledge occurs after editing. To alleviate this serious problem, we propose a simple editing framework METO (Multi-Editing with Time Objective) that can be applied to enhance existing editing methods easily.
@@
For a chain of temporal facts $\mathcal{C}=\{(s,r,o_{1},t_{s_{1}},t_{u_{1}}),...,(s,r,o_{N},t_{s_{N}},t_{u_{N}})\}$, the object is constantly changing and the knowledge of the model about $(s,r,\cdot)$ is determined by the collection time of its training corpus. Therefore, in order to better preserve the historical knowledge inside the model, we need to extract the model-time knowledge about $(s,r,\cdot)$. With the method in Section Locating Time of Model Facts, we obtain the fact of what is happening under the model time which is noted as $C_{m}=\{(s,r,o_{i},t_{s_{i}},t_{u_{i}})\}$. By comparing $\mathcal{C}$ and $C_{m}$, we can get the knowledge that needs to be newly captured by the model $C_{m+}=\{(s,r,o_{i+1},t_{s_{i+1}},t_{u_{i+1}}),...,(s,r,o_{N},t_{s_{N}},t_{u_{N}})\}$.
@@
Results on AToKe-SE and AToKe-ME